Learning Goal:

To practice Problem-Solving Strategy 8.1 Static equilibrium problems.

When you lift an object by moving only your forearm, the main lifting muscle in your arm is the biceps. Suppose the mass of a forearm is 1.40 kg . If the biceps is connected to the forearm a distance dbiceps = 3.50 cm from the elbow, how much force Fbiceps must the biceps exert to hold a 950 g ball at the end of the forearm at distance dball = 35.0 cm from the elbow, with the forearm parallel to the floor? How much force Felbow must the elbow exert?

What is the horizontal distance dforearm between the elbow and the point where the weight of the forearm acts?

Express your answer in meters.
Answer #1

A)

Taking moment about the elbow,

(0.95 x 9.8 x 0.35) + (1.4 x 9.8 x 0.175) = Fb (0.035)

Fb = 161.7 N

Fe = Fb - (0.95 + 1.4)(9.8) = 138.67 N

B)

Distance, d = 0.35/2 = 0.175 m

Comment in case any doubt please rate my answer.....